# Deep Ensembles for Epistemic Uncertainty: A Frequentist Perspective

### v2

@@ 3.1 Problem Setup
$\mathrm{I}\left(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n}\right)$
$=$ $\mathrm{H}\left(Y_{\mathrm{test}}|X_{\mathrm{test}},\mathcal{D}_{n}\right)-\mathrm{H}\left(Y_{\mathrm{test}}|\theta,X_{\mathrm{test}},\mathcal{D}_{n}\right)$
@@ 3.3 A Bootstrap Estimator
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n}):=$
$\mathrm{H}(\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})])-\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b}))].$
@@
We note that our theorem also holds if we use the classic multinomial weights for bootstrap instead of Dirichlet ones—see Appendix A.
Theorem 3.2 proves the asymptotic validity of the bootstrap estimator and provides a frequentist path to studying the Bayesian MI.
@@ 3.4 Distinguishing Algorithmic Randomness from Statistical Uncertainty
$=\mathrm{H}(\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]$
$+\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]-\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]$
$=\underbrace{\mathrm{H}(\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)]])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]}_{=:\mathrm{I}^{\textrm{resampling}}_{b}}$
$+\underbrace{\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}\left[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{s}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]\right]}_{=:\mathrm{I}^{\textrm{seeds}}_{b}}.$ (3)
@@
This is parallel to the term $\mathrm{I}^{\textrm{seeds}}_{b}$ above, but with the full data instead of a bootstrap subsample. Thus, the bootstrap estimator above can be viewed as approximately the MI from random seeds plus an additional term capturing variability in the dataset. In the experiments, we will see that the latter is often smaller than the former. In this way,
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})\approx\mathrm{I}^{\textrm{deep ensemble}}.$
This link gives a new frequentist motivation for deep ensembles. More precisely, since Sec. 3.3 proved that the bootstrap estimator is asymptotically correct for the true Bayesian target, the above relation indicates how deep ensembles succeed in capturing the Bayesian MI.
@@ 5 Related Work
5.1 Uncertainty Decomposition
MI as a measure of epistemic uncertainty has demonstrated success in active learning (Houlsby et al., 2011; Gal et al., 2017), out-of-distribution and misclassification detection tasks (Kotelevskii et al., 2025). Beyond the specific information theoretic uncertainty decomposition considered in our work, a variety of other ways of quantifying epistemic uncertainty have also been proposed (Senge et al., 2014; Kendall and Gal, 2017; Hofman et al., 2024; Lahlou et al., 2023; Schweighofer et al., 2023; Smith et al., 2025) and MI does have drawbacks (Wimmer et al., 2023). See (Hüllermeier and Waegeman, 2021) for a comprehensive review of the topic. Indeed, epistemic and aleatoric uncertainty are often not cleanly separable and take on different meanings in different contexts (Kirchhof et al., April 28, 2025). We use MI for our investigations because of its widespread use, strong conceptual foundation and practical success; our work is not intended to address the optimality of MI as an epistemic uncertainty assessment.
@@ 5.1 Uncertainty Decomposition
5.2 Bayesian and Frequentist Equivalence
(Rubin, 1981) considers bootstrap from a Bayesian lens, although this interpretation is hard to reconcile with traditional Bayesian inference since it considers priors over the data rather than the parameters. Closest to our theoretical results is the work by (Newton and Raftery, 1994) which studied the use of bootstrap as a way of generating samples from a Bayesian posterior. Our work expands upon this idea by specifically leveraging this connection to connect frequentist and Bayesian notions of uncertainty. Moreover, we operationalize this idea in a modern context with DNNs and use it to study deep ensembles. Finally, (Fong et al., 2023) also connects Bayesian parameter uncertainty to frequentist uncertainty arising due to lack of data. While their work conceptually shares connections to ours, it consider different practical applications.
5.3 Deep Ensembles
The uncertainty quantification capabilities of deep ensembles (Lakshminarayanan et al., 2017) were studied in depth by (Ovadia et al., 2019) and (Ashukha et al., 2020). Broadly speaking, these studies observed that deep ensembles have well calibrated probabilistic forecasts both in and out-of-domain, and exhibit strong performance on downstream tasks such as out-of-distribution detection and misclassification detection.
@@ 5.3 Deep Ensembles
5.4 Bootstrapping for Uncertainties in Deep Learning
The idea of using bootstrap to estimate uncertainties in machine learning has been explored by several works in different contexts. (Jürgens et al., 2024) investigates the suitability of evidential deep learning (Sensoy et al., 2018) for uncertainty quantification by comparing them to the bootstrapped estimates of the model parameters, which they believe accurately capture uncertainty in model predictions due to randomness in the training data. (Jiménez et al., 2025) investigates variance based measures of uncertainty in a regression setting to also note that deep ensembles only capture uncertainty due to training variability which often dominates over uncertainty due to randomness in training data, similar to our findings in Sec. 3. Lastly, (Shaker and Hüllermeier, 2020) have also considered approximating the posterior distribution by the trees in a random forest, which can be seen as bootstrap estimates, for the purposes of mutual information estimation as in our work.
@@ 6.2 Future Directions
Figure 4: Influence Function Approximation: Estimates of epistemic uncertainty using Dropout, versus using influence functions to approximate predictions from bootstrapped datasets.
@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@ 3.1 Problem Setup
$\mathrm{I}\left(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n}\right)$ $=\mathrm{H}\left(Y_{\mathrm{test}}|X_{\mathrm{test}},\mathcal{D}_{n}\right)-\mathrm{H}\left(Y_{\mathrm{test}}|\theta,X_{\mathrm{test}},\mathcal{D}_{n}\right)$
@@ 3.3 A Bootstrap Estimator
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n}):=\mathrm{H}(\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})])-\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b}))].$
@@
We note that our theorem also holds if we use the classic multinomial weights for bootstrap instead of Dirichlet ones—see Appendix A. Theorem 3.2 proves the asymptotic validity of the bootstrap estimator and provides a frequentist path to studying the Bayesian MI.
@@ 3.4 Distinguishing Algorithmic Randomness from Statistical Uncertainty
$=$ $\mathrm{H}(\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]+\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]-\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]$
$=$ $\underbrace{\mathrm{H}(\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)]])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]}_{=:\mathrm{I}^{\textrm{resampling}}_{b}}+\underbrace{\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}\left[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{s}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]\right]}_{=:\mathrm{I}^{\textrm{seeds}}_{b}}.$ (3)
@@
This is parallel to the term $\mathrm{I}^{\textrm{seeds}}_{b}$ above, but with the full data instead of a bootstrap subsample. Thus, the bootstrap estimator above can be viewed as approximately the MI from random seeds plus an additional term capturing variability in the dataset. In the experiments, we will see that the latter is often smaller than the former. In this way, $\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})\approx\mathrm{I}^{\textrm{deep ensemble}}.$ This link gives a new frequentist motivation for deep ensembles. More precisely, since Sec. 3.3 proved that the bootstrap estimator is asymptotically correct for the true Bayesian target, the above relation indicates how deep ensembles succeed in capturing the Bayesian MI.
@@ 4.4 Results
Figure 4: Influence Function Approximation: Estimates of epistemic uncertainty using Dropout, versus using influence functions to approximate predictions from bootstrapped datasets. See Sec. 6.2 for details.
@@ 5 Related Work
Uncertainty Decomposition: MI as a measure of epistemic uncertainty has demonstrated success in active learning (Houlsby et al., 2011; Gal et al., 2017), out-of-distribution and misclassification detection tasks (Kotelevskii et al., 2025). Beyond the specific information theoretic uncertainty decomposition considered in our work, a variety of other ways of quantifying epistemic uncertainty have also been proposed (Senge et al., 2014; Kendall and Gal, 2017; Hofman et al., 2024; Lahlou et al., 2023; Schweighofer et al., 2023; Smith et al., 2025) and MI does have drawbacks (Wimmer et al., 2023). See (Hüllermeier and Waegeman, 2021) for a comprehensive review of the topic. Indeed, epistemic and aleatoric uncertainty are often not cleanly separable and take on different meanings in different contexts (Kirchhof et al., April 28, 2025). We use MI for our investigations because of its widespread use, strong conceptual foundation and practical success; our work is not intended to address the optimality of MI as an epistemic uncertainty assessment.
@@
Bayesian and Frequentist Equivalence: (Rubin, 1981) considers bootstrap from a Bayesian lens, although this interpretation is hard to reconcile with traditional Bayesian inference since it considers priors over the data rather than the parameters. Closest to our theoretical results is the work by (Newton and Raftery, 1994) which studied the use of bootstrap as a way of generating samples from a Bayesian posterior. Our work expands upon this idea by specifically leveraging this connection to connect frequentist and Bayesian notions of uncertainty. Moreover, we operationalize this idea in a modern context with DNNs and use it to study deep ensembles. Finally, (Fong et al., 2023) also connects Bayesian parameter uncertainty to frequentist uncertainty arising due to lack of data. While their work conceptually shares connections to ours, it consider different practical applications.
Deep Ensembles: The uncertainty quantification capabilities of deep ensembles (Lakshminarayanan et al., 2017) were studied in depth by (Ovadia et al., 2019) and (Ashukha et al., 2020). Broadly speaking, these studies observed that deep ensembles have well calibrated probabilistic forecasts both in and out-of-domain, and exhibit strong performance on downstream tasks such as out-of-distribution detection and misclassification detection.
@@
Bootstrapping for Uncertainties in Deep Learning: The idea of using bootstrap to estimate uncertainties in machine learning has been explored by several works in different contexts. (Jürgens et al., 2024) investigates the suitability of evidential deep learning (Sensoy et al., 2018) for uncertainty quantification by comparing them to the bootstrapped estimates of the model parameters, which they believe accurately capture uncertainty in model predictions due to randomness in the training data. (Jiménez et al., 2025) investigates variance based measures of uncertainty in a regression setting to also note that deep ensembles only capture uncertainty due to training variability which often dominates over uncertainty due to randomness in training data, similar to our findings in Sec. 4. Lastly, (Shaker and Hüllermeier, 2020) have also considered approximating the posterior distribution by the trees in a random forest, which can be seen as bootstrap estimates, for the purposes of mutual information estimation as in our work.
@@ 7 Conclusion
Acknowledgments
The authors thank Willem Waegeman, Mira Jürgens and Sebastián Jiménez for their insightful comments and pointers to relevant literature. The authors would also like to thank Dan Kluger for his kind help and insights into the proofs. A.J. and S.B. were partly supported by the ARPA-H ADAPT program.